OpenStudy (nerdgamer_):

Medal if you help me. How would I explain the steps you would use to compare four-fifteenths and five-sixths.

OpenStudy (jdoe0001):

well.... one easy way to do such is make both the same denominator how? well say multiplying each fraction by the OTHER'S denominator \(\bf \cfrac{4}{\color{blue}{15}}\times\color{red}{ 6}\qquad \cfrac{5}{\color{red}{6}}\times \color{blue}{15}\)

OpenStudy (jdoe0001):

hmmm ahemm... lemme redo that...

OpenStudy (jdoe0001):

how? well say multiplying each fraction by the OTHER'S denominator \(\bf \cfrac{4}{15}\times \cfrac{6}{6}\qquad \cfrac{5}{6}\times \cfrac{15}{15}\)

OpenStudy (jdoe0001):

like so, once you do that, you'd end up with the same proportional denominator

OpenStudy (jdoe0001):

well. if you multiply those 2 fractions, what do you get for both?

OpenStudy (jdoe0001):

well, can't write what you wouldn't understand